OpenStudy (anonymous):

solve the differential equation: dy/dx = x+1/(x^2+2x-3)^2

OpenStudy (anonymous):

You isolate the x's with the dx and the dy on it's own on their own side. Doing this you multiply by dx to both right hand side and left hand side. Then you have dy=x+1/(x^2+2x-3)^2 dx. And to solve for this you integrate both sides.

OpenStudy (anonymous):

well its just integration with respect to x

OpenStudy (anonymous):

\[\int\limits_{}^{}dy=\int\limits_{}^{}(x+1)/(x^2+2x-3)^2dx\] \[y=\int\limits\limits_{}^{}(x+1)/(x^2+2x-3)^2dx\] Simplify the integrand on the left hand side and integrate :-)

OpenStudy (anonymous):

*integrand in the right hand side sorry LOL

OpenStudy (anonymous):

i dont know how to integrate this one... y= (x+1)(x^2+2x-3)^-2 ? do i multiply it out first and then integrate?

OpenStudy (anonymous):

Well actually I didn't even check to see if you can simplify it but let me try to work it out :-) Then I'll post.

OpenStudy (anonymous):

ok great thank you

OpenStudy (anonymous):

it is an inspection/substitution method do you see if we differentiate x^2+2x-3 we get double this: x+1

OpenStudy (anonymous):

Yes you have \[(x+1)/(x-1)^2(x+3)^2\] in the integrand

OpenStudy (anonymous):

so if we were to differentiate a function f(x^2+2x-3) we would get 2f'(x^2+2x-3)(x+1)

OpenStudy (anonymous):

alternatively use the substitution u = x^2+2x-3

OpenStudy (anonymous):

so we have\[\frac{du}{dx} = 2(x+1) \text{ , } dx = \frac{du}{2(x+1)}\] \[\int\limits{\frac{x+1}{(x^2+2x-3)^2}}dx = \int\limits{\frac{1}{2u^2}du}\]

OpenStudy (anonymous):

how did you get the 2 for 2u^2?

OpenStudy (anonymous):

where did the 1/2 come from?

OpenStudy (anonymous):

u was chosen to be \[u = x^2+2x-3\] so du=\[du=(2x+2)dx\] Solving for \[dx=du/(2x+2)=du/2(x+1) \] so when you put in back into the integrand you have the 1/2 to account for the 2 in the du equation

OpenStudy (anonymous):

krystabell is spot on :)

OpenStudy (anonymous):

ok thank you guys!